OpenStudy (alli6106):

if to the numerator and denominator of the fraction 1/3 you add its denominator, 3, the fraction will double [(1+3) / (3+3) = 4/6 = 2/3]. Find a fraction which will triple when its denominator is added to its numerator and denominator; find one that will quadruple.

OpenStudy (alli6106):

kind of

myininaya (myininaya):

what is it

myininaya (myininaya):

don't be afraid to share

myininaya (myininaya):

i'm still thinking as well

OpenStudy (alli6106):

It is really hard for me to explain

myininaya (myininaya):

So we have the equality \[\frac{a+b}{2b}=\frac{3a}{b}\]

myininaya (myininaya):

maybe we can play with this somehow

myininaya (myininaya):

\[\frac{\frac{a+b}{2}}{b}=\frac{3a}{b} \\ \frac{a+b}{2}=3a \\ a+b=6a \\ b=5a\]

myininaya (myininaya):

So maybe if we choose a a and a b such that b=5a and it might work...

myininaya (myininaya):

choose a random value a not equal to 0

OpenStudy (alli6106):

3

myininaya (myininaya):

ok so b=5a so b has to be what if a=3?

myininaya (myininaya):

b=5a b=5(3) <---because you chose a to be 3

myininaya (myininaya):

so b=?

OpenStudy (alli6106):

ummmm 125?

myininaya (myininaya):

3*5=15 \[\frac{a}{b}=\frac{3}{15}\]

OpenStudy (alli6106):

oh

myininaya (myininaya):

now see if that fraction works

OpenStudy (alli6106):

i kinda get it

myininaya (myininaya):

3/15 can actually be reduced before you begin

OpenStudy (alli6106):

to 1/5

myininaya (myininaya):

right now see what happens if you add 5 to top and bottom

myininaya (myininaya):

we want triple of what we started with

OpenStudy (alli6106):

15?

myininaya (myininaya):

have you added 5 to the top and added 5 to the bottom yet?

OpenStudy (alli6106):

or 25?

myininaya (myininaya):

\[\frac{1}{5} -> \frac{1+5}{5+5}=\frac{?}{?}\]

OpenStudy (alli6106):

6/10

OpenStudy (alli6106):

OR 2/3

OpenStudy (alli6106):

I MEAN 3/5

myininaya (myininaya):

guess what?

myininaya (myininaya):

isn't that triple what we started with?

myininaya (myininaya):

3*1/5=3/5

OpenStudy (alli6106):

YA

OpenStudy (alli6106):

THE ANSWER IS 3/5

myininaya (myininaya):

now all i did was try to get a relationship between a and b I could work with we had the equality \[\frac{a+b}{2b}=\frac{3a}{b} \\ \text{ then I played with the fraction on the left to get if over b } \\ \text{ so I could just set the tops equal } \\ \frac{\frac{a+b}{2}}{b}=\frac{3a}{b} \\ \\ \text{ so then we had } \frac{a+b}{2}=3a \\\text{ multiply } 3 \text{ on both sides } \\ a+b=6a \\ \text{ subtract } a \text{ on both sides } b=5a \] the fraction you started with was the answer we were checking to see if it was the answer by going through the steps your question had put forth and in genernal where a and b do not equal 0 we have \[\frac{a}{b}=\frac{a}{5a} \text{ will always work } \\ \frac{a}{5a} \text{ add 5a on top and bottom } \\ \frac{a+5a}{5a+5a}=\frac{6a}{10a}=3 \frac{2a}{10a}=3 \frac{a}{5a}\]

OpenStudy (alli6106):

OK THANK YOU

myininaya (myininaya):

you remember the fraction we started with was the one you made

OpenStudy (alli6106):

ok

myininaya (myininaya):

now see if you can come up with an answer for the next one

myininaya (myininaya):

"find one that will quadruple"

myininaya (myininaya):

also one more thing a/(5a)=1/5 for any value a not equal to zero :)

myininaya (myininaya):

you can use the same process i used above for the next question

myininaya (myininaya):

except with a few tweaks

OpenStudy (alli6106):

ok thank you